# Gravitational Wave Signatures from Periodic Orbits around a non-commutative inspired black hole surrounded by quintessence

### v1

@@ IV Gravitational Radiation in non-commutative geometry
To enhance the visual clarity of the plots for the characteristic strain given in Eq. (4.15), we applied a smoothing procedure to the numerically generated $h_{c}(f)$ by performing a running average over 30 frequency bins. Choosing a larger averaging window would further repress numerical noise but could also obscure fine spectral features. As shown in Fig. 7, portions of the characteristic strain corresponding to different orbital configurations $(z,w,v)$ and values of the noncommutative parameter $\Theta$ lie above the sensitivity curve of the Laser Interferometer Space Antenna (LISA). This means that the corresponding gravitational waves, exhibiting distinctive zoom–whirl features arising from spacetime noncommutativity, fall within the detectable range of LISA LISA:2017pwj; Gair:2017ynp; Babak:2017tow. Such detections would provide an important opportunity to probe the geometry of spacetime around supermassive black holes and test possible quantum gravity effects through precise GW observations.
@@ V Discussions and Conclusions
In particular, we investigate periodic orbits and their corresponding waveforms within the context of non-commutative geometry. From the geodesic equations for black holes, we can analytically solve the equations. Then we use a special taxonomy Levin:2008mq to distinguish different types of periodic orbits in non-commutative geometry. In this scheme, each periodic orbit is described by a set of parameters $(z,w,v)$. This study examined the effect of the non-commutative parameter $\Theta$ on the orbits of particles around a black hole. Unlike the classic Schwarzschild case, the presence of $\Theta$ significantly altered these orbits.
@@
In future work, we plan to extend this analysis in several ways. One important step will be to include the radiation reaction effects to study how gravitational wave emission changes the orbits over time. It will also be beneficial to extend the quadrupole approximation and incorporate higher multipole moments to achieve more accurate waveforms. Another natural extension is to consider rotating noncommutative black holes, where spin may further affect the orbital motion and emitted radiation. Finally, with future space-based detectors like LISA LISA:2017pwj and Taiji Ruan:2018tsw, improved waveform templates from such studies could help test deviations from general relativity and explore possible signatures of quantum gravity in the strong-field region. Finally, once these accurate waveforms become available, we will be able to investigate how future gravitational wave detectors might constrain or test the effects of dark matter on periodic orbits. We hope to address these challenges in future studies.

### v2

@@ IV Gravitational Radiation in non-commutative geometry
To enhance the visual clarity of the plots for the characteristic strain given in Eq. (4.15), we applied a smoothing procedure to the numerically generated $h_{c}(f)$ by performing a running average over 30 frequency bins. Choosing a larger averaging window would further repress numerical noise but could also obscure fine spectral features. As shown in Fig. 7, portions of the characteristic strain corresponding to different orbital configurations $(z,w,v)$ and values of the noncommutative parameter $\Theta$ lie above the sensitivity curve of the Laser Interferometer Space Antenna (LISA). This means that the corresponding gravitational waves, exhibiting distinctive zoom–whirl features arising from spacetime noncommutativity, fall within the detectable range of LISA Amaro-Seoane and others (2017); Gair et al. (2017); Babak et al. (2017). Such detections would provide an important opportunity to probe the geometry of spacetime around supermassive black holes and test possible quantum gravity effects through precise GW observations.
@@ V Discussions and Conclusions
In particular, we investigate periodic orbits and their corresponding waveforms within the context of non-commutative geometry. From the geodesic equations for black holes, we can analytically solve the equations. Then we use a special taxonomy Levin and Perez-Giz (2008) to distinguish different types of periodic orbits in non-commutative geometry. In this scheme, each periodic orbit is described by a set of parameters $(z,w,v)$. This study examined the effect of the non-commutative parameter $\Theta$ on the orbits of particles around a black hole. Unlike the classic Schwarzschild case, the presence of $\Theta$ significantly altered these orbits.
@@
In future work, we plan to extend this analysis in several ways. One important step will be to include the radiation reaction effects to study how gravitational wave emission changes the orbits over time. It will also be beneficial to extend the quadrupole approximation and incorporate higher multipole moments to achieve more accurate waveforms. Another natural extension is to consider rotating noncommutative black holes, where spin may further affect the orbital motion and emitted radiation. Finally, with future space-based detectors like LISA Amaro-Seoane and others (2017) and Taiji Ruan et al. (2020), improved waveform templates from such studies could help test deviations from general relativity and explore possible signatures of quantum gravity in the strong-field region. Finally, once these accurate waveforms become available, we will be able to investigate how future gravitational wave detectors might constrain or test the effects of dark matter on periodic orbits. We hope to address these challenges in future studies.